Circular No. 7670

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001dg IN CGCG 513-10 G. M. Hurst, Basingstoke, England, reports the discovery by M. Arm-

strong (Rolvenden, Kent) of a supernova (mag ∼ 17.5) on an unfiltered CCD image taken on July 22.023 UT in the course of the U.K. Nova/Supernova Patrol. SN 2001dg is located at α = 22^{h}10^{m}13^{s}.28, δ = +36^{o}39′14′′.2 (equinox 2000.0), as determined from a follow-up image obtained on July 23.024, when the new object appeared at similar brightness; this is 19′′ west and 3′′ south of the center of CGCG 513-10. No stellar object appears at this location on Armstrong’s image of the galaxy from 2000 Sept. 23 (limiting mag ∼ 19.0). R. Arbour (South Wonston, Hampshire) confirmed the presence of the new object on an exposure taken on July 22.908.

T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum of SN 2001dg, obtained by P. Berlind on July 24.40 UT with the Mt. Hopkins 1.5-m telescope (+ FAST spectrograph), shows it to be a type-II supernova near maximum light. The spectrum exhibits a blue continuum with P-Cyg lines of hydrogen. Adopting the NASA/IPAC Extragalactic Database recession velocity of 6143 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hβ line is ∼ 8600 km/s.

SUPERNOVA 2001dh IN MCG –06-44-26 R. Chassagne, Ste. Clotilde, Île de Réunion, reports the discovery of

an apparent supernova (mag ∼ 15.5) on unfiltered CCD images taken on July 22.88 and 24.85 UT with a 0.3-m reflector. The new object is located at α = 20^{h}17^{m}24^{s}.16, δ = −38^{o}40′37′′.4 (equinox 2000.0), which is 38′′.7 east and 11′′.2 south of the nucleus of MCG –06-44-26 (= ESO 340-G9). Images taken on June 18.1 and 26.1 show nothing at this position (limiting mag ∼ 19.0), and SN 2001dh does not appear on a Digital Sky Survey image taken with the U.K. Schmidt 1.2-m telescope on 1978 Sept. 24.4 (limiting mag ∼ 21.0).

WZ SAGITTAE “Quick-look” magnitudes (accurate to ∼ 2 percent) obtained by A. U.

Landolt, Louisiana State University, with the Cerro Tololo 1.5-m telescope: July 25.17728 UT, V = 8.26, B − V = −0.12, U − B = −0.99; 25.18021, 8.21, −0.11, −0.97, V − R = −0.02, R − I = −0.02; 25.23370, 8.31, −0.10, −0.96, −0.02, −0.04.

2001 July 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green